OpenStudy (anonymous):

write y - 9= -(x+2) in slope- intercept form.

OpenStudy (binary3i):

y= -x +7

OpenStudy (lgbasallote):

careful...

OpenStudy (lgbasallote):

i mean to the asker..not you:p

OpenStudy (lgbasallote):

dont forget to distribute the negative sign.. y - 9 = -x - 2 just a reminder ^_^

OpenStudy (binary3i):

yes @lgbasallote .